OpenStudy (anonymous):

what is the equation for a line that has a slope of -4/5 and contains point 0,1

OpenStudy (anonymous):

use the point slope formua.

OpenStudy (anonymous):

y-1=-4/5(x-0)

OpenStudy (anonymous):

y=-4/5x+1

OpenStudy (anonymous):

that's your answer

OpenStudy (anonymous):

slope point formula: (y-b)/(x-a)=m, where m is the slope and (a,b) is the point...use this and u'll get the answer as y=-4/5x+1